OpenStudy (anonymous):

in the measure of each side of a square is doubled, its area increases how many times?

OpenStudy (anonymous):

4

OpenStudy (aripotta):

4 times. you square the amount you changed the dimensions. for example, doubling each side: 2^2 = 4. so the area is 4 times the original also, tripling each side: 3^2 = 9. so the area would be 9 times the original

OpenStudy (anonymous):

Let's look at the equation of a square: \[area = side^2\] If we double a side, then we get: \[area = 2side^2\] 2^2 is equal to 4, so the area will increase by a factor of 4. Think about the picture of what will happen as well: